Example 8.8 Quantitative Calculations . A 200.0 mL sample of water was filtered through a preweighed glass fiber filter. After drying to constant weight at 105°C, the filter was found to have increased in mass by 48.2mg. Determine the total suspended solids for the sample in parts per million
Answer #1

Given data

Dry mass in 200 mL of water sample or suspended solids = 48.2 mg

Answer

Based on given data, dry mass in 1000 mL of water sample or suspended solids = 48.2 mg * 5 = 241 mg

Since 1000 ml = 1 L

Dry mass in 1 L of water sample or suspended solids = 241 mg

Dry mass in 1 L of water sample or suspended solids = 241 mg * (1*10-3) g /mg = 241 * 10-3 g

Since 1 L = 1000 cc

Dry mass in 1000 cc of water sample or suspended solids = 241 * 10-3 g

hence dry mass in 1 cc of water sample or suspended solids = 241 * 10-3 g / 1 * 103 cc = 241 * 10-6 g / cc ... eqn (1)

Now parts per million is the gram of suspended solids per million parts of water. Hence ppm = 1 * 10-6 g / cc... eqn (2)

From eqn (1) & eqn (2) the concentration of suspended solids = 241 ppm